OpenStudy (anonymous):

A balloon is being expanded such that its diameter d is increasing at a rate given by the function d = 1 + 3t, where d is in centimeters and t is the time in seconds. What is the volume of the balloon in cubic centimeters at time t? (Volume of a sphere is π6d3.)

OpenStudy (anonymous):

π(1 + 3t)2 (π3)(1+3t)2 (π6)(1+3t)3 π(1 + 3t)3

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

aintis c?

OpenStudy (anonymous):

I think its c

OpenStudy (freckles):

I don't understand there formula for a sphere but I see it is given as π6d3 whatever that means Like my volume of a sphere is \[\frac{4}{3} \pi r^3 \\ \frac{4 }{3} \pi (\frac{d}{2})^3 \\ \frac{4}{3} \pi \frac{d^3}{8} \\ \frac{\pi d^3}{6}\] and you are also given d=1+3t so yeah just replace the d above with (1+3t)

OpenStudy (anonymous):

so its C right?

OpenStudy (anonymous):

It is c